ID |
Date |
Author |
Type |
Category |
Subject |
5227
|
Sun Aug 14 00:26:51 2011 |
Nicole | Summary | SUS | TT Optimization Curves |
Quote: |
1) Drawing has the dimensions => You can calculate the volume => You can calculate the mass
Complicated structure can be ignored. We need a rough estimation.
2) Your restoring force can have two terms:
- one comes from the spring constant k
- the other from the gravity
|
Thank you.
The wire used to suspend the EDCs is tungsten?
To verify, for my model, the EDC will be the mass of all four dampers or a single damper? The length of the wire used to suspend the EDC will be the combined length of 4 wires or length of a single wire?
Taking into account the densities for each material (specific material of each component was listed, so I looked up the densities), and trying my best to approximate the volumes of each component, I have determined
the mass of the mirror + mirror holder to be ~100 g and the mass of a single EDC to be ~19 g |
5226
|
Sat Aug 13 21:48:17 2011 |
Koji | Summary | SUS | TT Optimization Curves |
1) Drawing has the dimensions => You can calculate the volume => You can calculate the mass
Complicated structure can be ignored. We need a rough estimation.
2) Your restoring force can have two terms:
- one comes from the spring constant k
- the other from the gravity
|
5225
|
Sat Aug 13 21:15:47 2011 |
Nicole | Summary | SUS | TT Optimization Curves |
Quote: |
What are the parameters you are using? As you have the drawings of the components, you can calculate the masses of the objects.
Reducing the ECD resonance from 10Hz->6Hz looks nice.
The resonant freq of the ECDs are not (fully) determined by the gravitational energy but have the contribution of the elastic energy of the wire.
Q1: How much is the res freq of the ECDs if the freq is completely determined by the grav energy? (i.e. the case of using much thinner wires)
Q2: How thin should the wires be?
|
The drawings do not have the masses of the objects.
For the resonant frequency:
Instead of sqrt (g/l) would the numerator in the square root be[ g + (energy stored in wire)/(mass of damper)] ?
|
5224
|
Sat Aug 13 19:08:01 2011 |
Koji | Summary | SUS | TT Optimization Curves |
What are the parameters you are using? As you have the drawings of the components, you can calculate the masses of the objects.
Reducing the ECD resonance from 10Hz->6Hz looks nice.
The resonant freq of the ECDs are not (fully) determined by the gravitational energy but have the contribution of the elastic energy of the wire.
Q1: How much is the res freq of the ECDs if the freq is completely determined by the grav energy? (i.e. the case of using much thinner wires)
Q2: How thin should the wires be? |
5223
|
Sat Aug 13 15:47:47 2011 |
Nicole | Summary | SUS | TT Optimization Curves |
Using my Matlab model of the flexibly-supported eddy current damping system, I have changed parameters to see if/how the TTs can be optimized in isolation. As I found earlier, posted in my bode plot entry, there is only a limited region where the flexibly-supported system provides better isolation than the rigidly-supported system.
Here is what I have found, where \gamma is the scale factor of the magnetic strength (proportional to magnetic strength), \beta is the scale factor of the current damper mass (estimated by attempting to fit my model to the experimental data), and \alpha is the scale factor of the current resonant frequency of the dampers.
 
Here are my commentaries on these plots. If you have any commentaries, it would be very helpful, as I would like to incorporate this information in my powerpoint presentation.
It seems as if the TT suspensions are already optimized?
It may be difficult to lower the resonant frequency of the dampers because that would mean changing the lengths of the EDC suspensions). Also, it appears that a rather drastic reduction (at most 0.6*current EDC resonant frequency --> reduction from about 10 Hz to 6 Hz or less) is required . Using the calculation that the resonant frequency is sqrt(g/length), for my single-suspended EDC model, this means increasing the wire length to nearly 3 x its current value. I'm not sure how this would translate to four EDCs...
The amplification at resonance caused by increasing the magnet strength almost offsets the isolation benefits of increasing magnet strength. From my modeling, it appears that the magnet strength may be very close (if not already at) isolation optimization.
Lowering the mass to 0.2 the current mass may be impractical. It seems as if the benefits of lowering the mass only occur when the mass is reduced by a factor of 0.2 (maybe 0.4)
|
5222
|
Sat Aug 13 15:40:38 2011 |
Nicole | Summary | SUS | TT Shaking Today and Hopefully More? |
As reported in my previous entry of TT supsension bode plots, I found that my experimental data had what appears to be very noise peaks above 20 Hz (as mentioned earlier, the peak at 22 Hz is likely due to vertical coupling, as 22 Hz is the resonant frequency of the cantilever blades). This is very unusual and needs to be explored further. I would like to vertically-shake the TTs to obtain more data on possible coupling. However, I am leaving on Monday and will not return until Thursday (day of SURF talks). I am leaving campus Friday afternoon or so. I would may need some help coming up with an assembly plan/assembling set-up for vertical shaking (if it is possible to do so in such a limited time frame).
Today I wanted to see if the "noisy peaks" above 30 Hz were due to EM noise coupling. I tested this hypothesis today, seeing if EM fields generated by the coil at higher frequencies were injecting noise into my transfer function measurements. I found that the "noisy peaks" above 30 Hz are NOT DUE TO EM NOISE COUPLING. I am very curious as to what is causing the high peaks (possibly coupling from other degrees of freedom)?
 |
5221
|
Sat Aug 13 02:31:42 2011 |
kiwamu | Update | SUS | Re: ETMY hopefully good again |
I guess the ETMY suspension is still fine. Their OSEM DC voltage and the free swinging spectra look healthy.
It could be a failure in the initial guess for fitting.
Quote from #5216 |
I'm no longer convinced that ETMY is healthy. I can't fit the peaks to get the input matrix.
|
|
5220
|
Sat Aug 13 02:11:33 2011 |
kiwamu | Update | SUS | free swinging again |
I am leaving all of the suspensions free swinging. They will automatically recover after 5 hours from now.
--
Excited all optics
Sat Aug 13 02:08:07 PDT 2011
997261703
--
FYI : I ran a combination of two scripts: ./freeswing && ./opticshutdown |
5219
|
Sat Aug 13 01:54:18 2011 |
kiwamu | Update | SUS | PRM OSEM adjustment part II |
Adjustment of the PRM OSEMs are done. The coils turned out to be healthy.
The malfunction was fixed. It was because the UL OSEM was too deeply inserted and barely touching the AR surface of the mirror.
(OSEM adjustment)
+ Excited POS at 6.5 Hz with an amplitude of 3000 cnts by the LOCKIN oscillator.
+ Looked at the signal of each sensor in frequency domain.
+ Maximized the excitation peak for each of the four face OSEMs by rotating them.
+ Minimized the excitation peak in the SIDE signal by rotating it.
+ Adjusted the OSEM translational position so that they are in the midpoint of the OSEM range.
(POS sensitivity check)
From the view point of the matrix inversion, one thing we want to have is the equally sensitive face sensors and insensitive SIDE OSEM to the POS motion.
To check the success level of today's PRM adjusment, I ran swept sine measurements to take the transfer function from POS to each sensor.
The plots below are the results. The first figure is the one measured before the adjustment and the second plot is the one after the adjustment.
As shown in the plot, before the adjustment the sensitivity of OSEMs were very different and the SIDE OSEM is quite sensitive to the POS motion.
So PRM used be in an extremely bad situation.
After the adjustment, the plot became much better.
The four face sensors have almost the same sensitivity (within factor of 3) and the SIDE is quite insensitive to the POS motion.

Quote from #5203 |
We tried adjusting the OSEMs on PRM, but we didn't complete it due to a malfunction on the coils.
The UL and LL coils are not working correctly, the forces are weak.
|
|
5218
|
Sat Aug 13 01:52:07 2011 |
Yoichi | Update | LSC | Feed forward delay |
Yoichi, Koji
While I was testing the feed forward cancellation, I noticed that the
cancellation was not perfect.
The test I did was the following.
I injected the same signal to both DARM and MICH feedback filters.
This was done by injecting a signal into the excitation point of
the ASDC PD, then changing the input matrix elements so that the signal
goes to both DARM and MICH.
Then in the FFC, MICH signal was fed forward to DARM by the gain of -1.
Ideally, this should completely eliminate the DARM FB signal.
In reality, it did not.
The first PDF compares the spectrum of the injected noise (white noise,
red curve) with the spectrum of the signal after the FFC (blue curve).
At higher frequencies, the cancellation becomes poor.
It suggests that this is caused by some delay in the FFC.
I also took a transfer function from the injection point to the signal
after the FFC (second attachment).
I fitted the measured TF with a theoretical formula of
1-exp(-i*dt*f),
where dt is the time delay and f is the frequency.
The fitting is very good, and I got dt = 0.8msec ~ 13 samples for 16kHz.
13 samples is something very large.
The cause of the delay was suspected to be the shared memory communication
between different processes.
I moved all the FFC blocks to c1lsc.mdl.
Then the cancellation becomes perfect. The signal after the FFC is
completely zero, so I couldn't even make a TF measurement.
This results suggest that a large delay of 13 samples is induced
when you use shared memory to send signals round trip.
We should make simpler models, just passing signals back and forth
via shared memory, dolphin network or GE FANAC RFM to check the
delays more precisely.
For the moment, the FCC is included in the c1lsc model.
The MEDM screens were modified to account for this change.
c1ffc is stopped and removed from rtsystab. |
5217
|
Fri Aug 12 20:33:57 2011 |
Dmass | Summary | PSL | NPRO PDH-Locked to Ref Cav |
To aid Jenny's valiant attempt to finish her SURF project, I did some things with the front end system over the last couple days, largely tricking Jamie into doing things for me lest I ruin the 40m RCG system. Several tribulations have been omitted.
We stole a channel in the frontend, in the proccess:
- Modified the C1GFD simulink model (now analog) to be "ADC -> TMP -> DAC" where TMP is a filter bank
- C1GFD_TMP.adl (in /opt/rtcds/caltech/c1/medm/c1gfd) is the relevant part which connects the ADC to the DAC in the frontend
- Confirmed that the ADC was working by putting a signal in and seeing it in the frontend
- Could not get a signal out of the anti aliasing board
- Looked sad until Kiwamu found a breakout board for the SCSI cable coming from the DAC
- Used SR560 to buffer DAC output
- drove a triangle wave with AWG into the TMP EXC channel (100 counts 1 Hz) and looked at it after the ~25 ft of BNC cable running between the DAC and the NRPO driver
- wave looked funny (not like a triangle wave), maybe the DAC is not meant to push a signal so far, so added buffer
- Took the control signal going to the fast input of the NPRO driver (using the 500 Ohm SR560 output - see Jenny's diagram) and put it into the anti aliasing board of the ADC
- Added switchable integrator to filter bank with Foton
- I couldn't get the names to display in the filter bank, so I looked sad again
- Jamie and Koji both poked at the "no name displayed" problem but had no conclusions, so I decided to ignore it
- I confirm that when the two filters were toggled "on" that the transfer function was as expected: simple integrator with a unity gain at ~10mHz - agrees with what Foton's Bode Plot tool says it should be (see attached DTT plot)
- I got Jamie to manually add the two epics channels from the TMP model to the appropriate .ini file so they would be recorded
- C1:GFD-TMP_OUTPUT (16 Hz)
- C1:GFD-TMP_INMON (16 Hz)
- RefCav heater servo seems to still be set up, so we can use existing channels:
- C1:PSL-FSS_RCPID_SETPOINT (temp setpoint - will do +/-1C steps about 35 C)
- C1:PSL-FSS_MINCOMEAS (In loop temp sensor - in C)
- C1:PSL-FSS_RCTEMP (out of loop temp sensor - in C)
- C1:PSL-FSS_TIDALSET (Voltage to heater - rails @ +/- 2V)
- Closed loop on the control signal for the NPRO driver with an integrator, saw error signal go to zero
- Turned up gain a little bit, saw some oscillations, then turned gain down to stop them, final gain = 2
- Left system on for Jenny to come in and do step responses
|
5216
|
Fri Aug 12 20:28:13 2011 |
Jenne | Update | SUS | ETMY hopefully good again |
Quote: |
[Jamie, Jenne]
We went in to have a look-see at ETMY since it looked stuck-ish. Jamie noticed that the side magnet was pretty close to the teflon plates of the OSEM. We rotated it a bit, and now its all better. We also adjusted the OSEMs until their mid-ranges were happy. The U's were a little low, and the L's were a little high, as if the optic were a bit pitched backward. Anyhow, we checked that the table is level, and tweaked the OSEMs. We're starting the free-swinging test now...
Excited all optics
Fri Aug 12 17:38:53 PDT 2011
997231148
|
Hmmm. I'm no longer convinced that ETMY is healthy. I think that when I gave it a kick, it's bouncing against something. I can't fit the peaks to get the input matrix. I guess step 1 is to try giving it a smaller kick for the free swinging spectra. But if the owl shift folk feel like it, they might have a look-see. |
5215
|
Fri Aug 12 17:37:11 2011 |
Jenne | Update | SUS | ETMY hopefully good again |
[Jamie, Jenne]
We went in to have a look-see at ETMY since it looked stuck-ish. Jamie noticed that the side magnet was pretty close to the teflon plates of the OSEM. We rotated it a bit, and now its all better. We also adjusted the OSEMs until their mid-ranges were happy. The U's were a little low, and the L's were a little high, as if the optic were a bit pitched backward. Anyhow, we checked that the table is level, and tweaked the OSEMs. We're starting the free-swinging test now...
Excited all optics
Fri Aug 12 17:38:53 PDT 2011
997231148 |
5214
|
Fri Aug 12 17:27:49 2011 |
Yoichi | Summary | CDS | Toggle button for RCG |
Bottom line: I made an RCG block to realize a toggle button easily.
Read on if you need such a button, or if you want to know how to
write a new RCG block with C.
-----------------
When I was making MEDM screens for FFC, I wanted to have a toggle
button to enable/disable the FFC path.
I wanted to have something like the ON/OFF buttons of the filter bank
screens, the one changes its state every time I click on it.
However, I could not find an easy way to realize that.
From MEDM, I can send a value to an EPICS channel using a "Message Button".
This value is always the same, say 1.
In the RCG model, I used a cdsEpicsMomentary block so that whenever the channel
gets 1, it stays to be 1 for a while and turns back to 0 in a second or so.
This generates a pulse of 1 when I click on a message button on a MEDM screen.
Then I needed a block to keep its internal state (0 or 1), and flips its state
whenever it receives a pulse of 1.
Since I couldn't find such a block in the current RCG library, I implemented one
using the cdsFunctionCall block. It allows you to implement a block with C code.
There is a good explanation of how to use this block in the CDS_PARTS library.
Here is basically what I did.
(1) Drag and drop thee cdsFunctionCall block to my model.
(2) In the "Block Properties", I put the following line in the Description field.
inline cdsToggle /opt/rtcds/caltech/c1/userapps/release/cds/common/src/cdsToggle.c
This means to call a function cdsToggle(), whose code is in the file indicated above.
(3) The contents of the source code is very simple.
void cdsToggle(double *in, int inSize, double *out, int outSize){
static double x = 0;
static double y = 0;
if (*in != y){
y = *in;
if (y == 1){
x = (x == 1) ? 0 : 1;
*out = x;
}
}
}
The function prototype is always the same. *in and *out are the pointers to the arrays of doubles
for input and output signals of the block. In simuLink, the signals have to be
multiplexed so that the RCG can know how many signals are handed to or returned from the function.
In order to keep the internal state of my custom block, I used "static" keyword in the
declaration of the variables. The rest of the code should be obvious.
(4) Just compile the model as usual. The RCG will automatically include the source code and put
a call to the function in the proper place.
I made the block a library so that people can use it.
/opt/rtcds/caltech/c1/userapps/trunk/cds/common/models/cdsToggle.mdl
is the one.
For the usage of it, please have a look at
/opt/rtcds/caltech/c1/userapps/trunk/isc/c1/models/c1lsc |
5213
|
Fri Aug 12 17:05:22 2011 |
Manuel, Ishwita | Configuration | PEM | STS2 Cable configuration |
The WWF_M connector is the end of the STS2 seismometer orange cable and the S1 connector is the end of the gray 26-pin-cable

|
5212
|
Fri Aug 12 16:52:49 2011 |
steve | Update | General | Chamber illuminator switch |
I'm looking for an ether net based power switch for turning lights on and off for the vacuum system from MEDM screen. This is what I found
Jamie please take a look at it.
|
5211
|
Fri Aug 12 16:50:37 2011 |
Yoichi | Configuration | CDS | FE Status screen rearranged |
I rearranged the FE_STATUS.adl so that I have a space to add c1ffc in the screen.
So, please be aware that the FE monitors are no longer in their original positions
in the screen. |
5210
|
Fri Aug 12 16:42:51 2011 |
Yoichi | Configuration | LSC | LSC Feed Forward Compensator |
I've been working on adding feed forward (FF) paths to our LSC code.
So far, I've made a basic feed forward functionality connected
to the feedback path of the LSC model.
As is shown in the MEDM screen, this feed forward compensator (FFC)
takes feedback signals from several DOFs (MICH, PRCL, SRCL, CARM, XARM and YARM)
and put those signals through some filters. Then the filtered signals are
summed into the feedback signals.
There are input and output matrices to select which signal goes to which signal.
Usually, we just want to feed forward MICH to DARM. We may also want to do PRCL
to DARM and SRCL to DARM if necessary.
It is more unlikely that we use CARM for FF. But I put it there just in case.
XARM and YARM will not going to be used as is. These are place holders for future
experiments, like low frequency FF from seismic channels or something like that.
Feed forward is almost always done to DARM. But just in case we want to do some
fancier FF, like FF from PRCL to MICH, the output matrix is there to chose where
the signals will go.
I haven't really tested it because we don't have the interferometer working.
But I checked the signal flow, and it seems the model is working correctly.
=== Implementation ===
FFC is running in a separate realtime code, called c1ffc.
This is to offload c1lsc from the possible intense calculations, like adaptive stuff,
performed in the FFC in the future.
The LSC signal is passed to c1ffc via shared memory. The calculated FF signals
are passed back to c1lsc via shared memory too.
Even though FFC is in a separate realtime model, it is still conceptually a part of LSC.
So, I used top_names tag to change the names of the channels to C1:LSC-FFC_* instead of
C1:FFC-*.
In MEDM, there is an "ENABLE" button in the FF screen. Even though it is shown in the FFC
overview screen, the button itself is in the c1lsc code, so that we can disable the FFC
even when c1ffc is dead or going crazy.
=== Background ideas ===
For those of you wondering what is this feed forward thing for, I will put a brief explanation here.
Taking MICH as an example, we get the error signal for MICH from probably REFL_55Q (or AS_55Q ?).
At low frequencies, this signal properly reflects the motion of the mirrors (mostly seismic).
However, it has much worse shot noise than DARM. At higher frequencies, like above a few tens of Hz,
the error signal is dominated by the shot noise. Feeding back this signal to, say BS, means
we are shaking the BS by the shot noise, which was otherwise quiet at high frequencies.
Now, if the BS is shaken, it has some intrinsic coupling to the DARM signal.
The mechanism is that the BS motion creates some audio frequency sidebands
and this SBs reach the AS port and beat against the local oscillator to create
fake GW signals. This is called "Loop Noise Coupling".
A well known way to mitigate this problem is feed forward.
Since we know how much we are shaking the BS (because we are doing it), and
we can measure the amount of BS to DARM coupling, we can subtract out the
loop noise by feeding forward the MICH feedback signal to the DARM actuators.
In other words, the noise SBs created at the BS is canceled on the PD by the
extra SBs created at the ETMs by the feed forward.
This is what FFC is trying to do. |
5209
|
Fri Aug 12 15:46:51 2011 |
Jenne | Update | SUS | Earthquake stop procedure |
According to Rana, the following is the "new" (should always have been used, but now we're going to enforce it) earthquake stop backing-off procedure:
1. Back all EQ stops away from the optic, so that it is fully free-swinging.
2. Confirm on dataviewer that the optic is truely free-swinging.
3. One at a time, slowly move the EQ stop in until it barely touches the optic. Watch dataviewer during this procedure - as soon as the time series of the OSEMs gets a 'kink', you've just barely touched the optic.
4. Back the EQ stop off by the calculated number of turns. No inspections, no creativity, just math. Each EQ stop should be between 1.5m and 2.0mm away from the optic.
5. Repeat steps 3 and 4 for each EQ stop.
Note: The amount that you need to turn the screws depends on what the threads are.
FACE and TOP stops are all 1/4-20, so 1.5 turns is 1.90mm
BOTTOM stops are either #4-40 or #6-32 (depending on the suspension tower). If #4-40, 3 turns is 1.90mm. If #6-32, 2.5 turns is 1.98mm |
5208
|
Fri Aug 12 15:34:16 2011 |
Nicole | Summary | SUS | Bode Plot for TT Suspension |
Quote: |
Here is my bode plot comparing the flexibly-supported and rigidly-supported EDCs (both with no bar)
It seems as if the rigidly-supported EDC has better isolation below 10 Hz (the mathematically-determined Matlab model predicted this...that for the same magnet strength, the rigid system would have a lower Q than the flexible system). Above 10 Hz (the resonance for the flexibly-supported EDCs seem to be at 9.8 Hz) , we can see that the flexibly-supported EDC has slightly better isolation? I may need to take additional measurements of the transfer function of the flexibly-supported EDC (20 Hz to 100 Hz?) to hopefully get a less-noisy transfer function at higher frequencies. The isolation does not appear to be that much better in the noisy region (above 20Hz). This may be because of the noise (possibly from the electromagnetic field from the shaker interfering with the magnets in the TT?). There is a 3rd resonance peak at about 22 Hz. I'm not sure what causes this peak...I want to confirm it with an FFT measurement of the flexibly-supported EDC (20 Hz to 40 Hz?)

|
Since the last post, I have found from the Characterization of TT data (from Jenne) that the resonant frequency of the cantilever springs for TT #4 (the model I am using) have a resonant frequency at 22 Hz. They are in fact inducing the 3rd resonance peak.
Here is a bode plot (CORRECTLY SCALED) comparing the rigidly-supported EDCs (model and experimental transfer functions)

Here is a bode plot comparing the flexibly-supported EDCs (model and experimental transfer functions). I have been working on this graph for FOREVER and with the set parameters, this is is close as I can get it (I've been mixing and matching parameters for well over an hour > <). I think that experimentally, the TTs have better isolation than the model because they have additional damping properties (i.e. cantilever blades that cause resonance peak at 22 Hz). Also, there may be a slight deviation because my model assumes that all four EDCs are a single EDC.

|
5207
|
Fri Aug 12 15:16:56 2011 |
jamie | Update | SUS | today's SUS overview |
Here's an update of the suspensions, after yesterdays in-vacuum work and OSEM tweaking:
- PRM and ETMY are completely messed up. The spectra are so bad that I'm not going to bother posting anything. ETMY has extreme sensor voltages that indicate that it's maybe stuck to one of the OSEMS. PRM voltages look nominal, so I have no idea what's going on there.
- ITMY is much improved, but it could still use some work
- SRM is a little worse than what it was yesterday, but we've done a lot of work on the ITMY table, such as moving ITMY suspension and rebalancing the table.
- BS looks for some reason slightly better than it did yesterday
|
5206
|
Fri Aug 12 14:15:07 2011 |
Nicole | Summary | SUS | Bode Plot for TT Suspension |
Here is my bode plot comparing the flexibly-supported and rigidly-supported EDCs (both with no bar)
It seems as if the rigidly-supported EDC has better isolation below 10 Hz (the mathematically-determined Matlab model predicted this...that for the same magnet strength, the rigid system would have a lower Q than the flexible system). Above 10 Hz (the resonance for the flexibly-supported EDCs seem to be at 9.8 Hz) , we can see that the flexibly-supported EDC has slightly better isolation? I may need to take additional measurements of the transfer function of the flexibly-supported EDC (20 Hz to 100 Hz?) to hopefully get a less-noisy transfer function at higher frequencies. The isolation does not appear to be that much better in the noisy region (above 20Hz). This may be because of the noise (possibly from the electromagnetic field from the shaker interfering with the magnets in the TT?). There is a 3rd resonance peak at about 22 Hz. I'm not sure what causes this peak...I want to confirm it with an FFT measurement of the flexibly-supported EDC (20 Hz to 40 Hz?)

|
5205
|
Fri Aug 12 11:07:50 2011 |
Nicole | Summary | SUS | More TT Shaking Completed This Morning |
This morning (about 10am to 11am), I have collected additional transfer function measurements for the T.T. suspension. I have finished taking my measurements. The SR785 has been returned to its place next the the seismometer racks.
The data has been backed up onto the cit40m computer |
5204
|
Fri Aug 12 04:11:13 2011 |
kiwamu | Update | SUS | free swinging |
Excited all optics - -
Fri Aug 12 03:34:12 PDT 2011
997180467
|
5203
|
Fri Aug 12 03:52:51 2011 |
kiwamu | Update | SUS | PRM OSEM adjustment |
[Suresh / Kiwamu]
We tried adjusting the OSEMs on PRM, but we didn't complete it due to a malfunction on the coils.
The UL and LL coils are not working correctly, the forces are weak.
Tomorrow we will look into the satellite box, which is one of the suspects.
During the adjustment we found that the POS excitation force was unequal in each sensor.
At the beginning we thought it's because of the difference of the sensitivity in each OSEM due to the bad OSEM orientations.
However it turned out that it comes from the actual force imbalance on each coil.
We checked the force of each coil by putting an offset (-2000 cnts) in each output digital filter and looked at the OSEM signals in time series.
The UL and LL coils are too weak and the responses are almost buried in the noise of the OSEMs in time series.
We briefly checked some analog electronics and found the DAC, AI board and deWhitening board were healthy.
We were able to see the right amount of voltage from the monitor pin on the front panel of the coil driver.
So something downstream are suspicious, including the satellite box, feedthrough and coils.
- - -
Although the coil issue, it could be worth trying to check the input matrix. |
5202
|
Fri Aug 12 03:49:45 2011 |
Jenny | Summary | PSL | NPRO PDH-Locked to Ref Cav |
DMass and I locked the NPRO laser (Model M126-1064-700, S/N 238) on the AP table to the reference cavity on the PSL table using the PDH locking setup shown in the block diagram below (the part with the blue background):

A Marconi IFR 2023A signal generator outputs a sine wave at 230 kHz and 13 dBm, which is split. One output of the splitter drives the laser PZT while the other is sent to a 7dBm mixer. Also sent to the mixer is the output of a photodiode that is detecting the reflected power from off the cavity. (A DC block is used so that only RF signal from the PD is sent to the mixer). The output of the mixer goes through an SR560 low-noise preamp, which is set to act as a low pass filter with a gain of 5 and a pole at 30 kHz. That error signal is then sent to the –B port of the LB1005 PDH servo, which has the following settings: PI corner at 10kHz, LF gain limit of 50 dB, and gain of 2.7 (1.74 corresponds to a decade, so the signal is multiplied by 35). The output signal from the LB1005 is added to the 230 kHz dither using another SR560 preamp, and the sum of the signals drive the PZT.
I am monitoring the transmission through the cavity on a digital oscilloscope (not shown in the diagram) and with a camera connected to a TV monitor. I sweep the NPRO laser temperature set point manually until the 0,0 mode of the carrier frequency resonates in the cavity and is visible on the monitor. Then I close the loop and turn on the integrator on the LB1005.
The laser locks to the cavity both when the error signal is sent into the A port and when it is sent into the –B port of the PDH servo. I determined that –B is the right sign by comparing the transmission through the cavity on the oscilloscope for both ways.
When using the A port, the transmission when it was locked swept from ~50 to ~200 mV (over ~10 second intervals) but had large high frequency fluctuations of around +/- 50 mV. Looking at the error signal on the oscilloscope as well, the RMS fluctuations of the error signal were at best ~40 mV peak to peak, which was at a gain of 2.9 on the LB1005.
Using the –B port yielded a transmission that swept from 50 to 250 mV but had smaller high frequency fluctuations of around +/- 20 mV. The error signal RMS was at best 10mV peak to peak, which was at a gain of 2.7. (Although over the course of 10 minutes the gain for which the error signal RMS was smallest would drift up or down by ~0.1).
The open loop error signal peak-to-peak voltage was 180 mV, which is more than an order of magnitude larger than the RMS error signal fluctuations when the loop is closed, indicating that it is staying in the range in which the response is linear.

In the above plot the transmission signal is offset by 0.1 V for clarity.
Below is the closed loop error signal. The inset plot shows the signal viewed over a 1.6 ms time period. You can see ~60 microsecond fluctuations in the signal (~17 kHz)

The system remained locked for ~45 minutes, and may have stayed locked for much longer, but I stopped it by opening the loop and turning off the function generator. Below is a picture of the transmitted light showing up on a monitor, the electronics I'm using, and a semi-ridiculous mess of wires.

I determined that it’s not dangerous to leave the system locked and leave for a while. The maximum voltage that the SR560 will output to the PZT is 10Vpp. This means that it will not drive the PZT at more than +/-5 V DC. At low modulation rates, the PZT can take a voltage on the order of 30 Vpp, according to the Lightwave Series 125-126 user’s manual, so the control signal will not push the PZT too hard such that it’s harmful to the laser.
|
5201
|
Fri Aug 12 00:18:30 2011 |
Ishwita, Manuel | Update | PEM | Coherence of Guralp1 and STS2(Bacardi, Serial NR 100151) |
We moved the seismometer STS2(Bacardi, Serial NR 100151) as we told in this Elog Entry, so the distance between Guralp1 and STS2 is 31.1m. Following is the coherence plot for this case:

then we also moved the Guralp1 under the BS and plugged it with the Guralp2 cable (at 7:35pm PDT), so now the distance between the two seismometers is 38.5m. Following is the coherence plot for this case:

|
5200
|
Thu Aug 11 19:14:22 2011 |
Ishwita , Manuel | Update | PEM | Seismometer STS2(Bacardi, Serial NR 100151) moved near ETMX |
We moved the STS2(Bacardi, Serial NR 100151) to his new location and laid his cable from rack 1X7 to ETMX. The seismometer was below the mode cleaner vacuum tube before.
Now, (since 6:05pm PDT) its placed near the ETMX.
 
|
5199
|
Thu Aug 11 19:01:45 2011 |
Suresh | Update | SUS | ITMX given a kick to start it freeswing' but to no avail. |
We started an ITMX freeswing run at this time
Thu Aug 11 18:58:59 PDT 2011
997149554
But the optic did not repond to the kick. It is possible that the earthquake stops are close to the face and/or rear of the optic and prevent it from oscillating. We will check again and see what is up in a few hours.
|
5198
|
Thu Aug 11 18:30:40 2011 |
Koji | Update | SUS | ITMY OSEM adjustment |
[Jamie, Koji]
ITMX OSEMs were adjusted so as to have the right DC numbers and the more uniform response to POS excitation.
It is waiting for the free-swinging test.
- ITMX was moved from its position to the north side of the table.
- The table was rebalanced.
- We found that the output of the LR OSEM has an excess noise compared with the other OSEMs.
We tried to swap the LR and SD OSEMs, but the SD OSEM (placed at the LR magnet) showed
the same excess noise at around 10-50Hz.
- We found that one of the EQ stops was touching the mirror. By removing this friction, all of the OSEMs
come to show similar power spectra. Good!
- Then we started to use LOCKIN technique to measure the sensitivity of the OSEMs to the POS excitation.
Originally the response of the OSEMs was as follows
UL 3.4 UR 4.3
LL 0 LR 2.5
After the adjustment of the DC values, final values became as follows
UL 3.9 UR 4.4
LL 3.9 LR 3.2
- We decided to close the light door. |
5197
|
Thu Aug 11 16:21:16 2011 |
Ishwita , Manuel | Update | PEM | Power spectra and Coherence of Guralps and STS2 |
Following is the power spectrum plot (with corrected calibration [see here]) of seismometers Guralp1 and STS2(Bacardi, Serial NR 100151):

The seismometers are placed approximately below the center of the mode cleaner vacuum tube. |
5196
|
Thu Aug 11 16:15:59 2011 |
Ishwita, Manuel | Update | PEM | Calibration of Guralp and STS2 |
Finally, we have found the correct calibration of Guralp and STS2 seismometers.
ADC: 216counts = 20V Hence, calibration of ADC is 3.2768e+03 counts/V.
GURALP
Sensitivity of seismometer = 800 V/ms-1
Gain of the guralp breakout box (reference elog entry) = 20
Calibration = 3.2768e+03 counts/V x 800 V/ms-1 x 20 = 52428800 counts/ms-1 -----> 1.9073e-08 ms-1/count
STS
Sensitivity = 1500 V/ms-1
Gain of the STS electronic breakout box = 10
Calibration = 3.2768e+03 counts/V x 1500 V/ms-1 x 10 = 49152000 counts/ms-1 -----> 2.0345e-08 ms-1/count
|
5195
|
Thu Aug 11 16:09:05 2011 |
Nicole | Summary | SUS | Bode Plot for TT Suspension |
All of my plots have already taken into account the calibration of the photosensor (V/mm ratio)
Here is a bode plot generated for the transfer function measurements we obtained last night/this morning. This is a bode plot for the fully-assembled T.T. (with flexibly-supported dampers and bottom bar). I will continue to upload bode plots (editing this post) as I finish them but for now I will go to sleep and come back later on today.

Here is a bode plot comparing the no eddy-current damper case with and without the bar that we suspected to induce some non-uniform damping. We have limited data on the NO EDC, no bar measurements (sine swept data from 7 Hz to 50 Hz) and FFT data from 0 Hz to 12.5 Hz because we did not want to induce too much movement in the mirror (didn't want to break the mirror). This plot shows that there is not much difference in the transfer functions of the TT (no EDC) with and without the bar.

From FFT measurements of the no eddy-current damper case without the bar (800 data points, integrated 10 times) we can define the resonance peak of the TT mirror (although there are still damping effects from the cantilever blades).
The largest resonance peak occurs at about 1.94 Hz. The response (magnitude) is 230.
The second-largest resonance peak occurs at about 1.67 Hz. The response (magnitude) is 153. This second resonance peak may be due to pitch motion coupling (this is caused by the fact that the clamping attaching the mirror to the wires occurs above the mirror's center of mass, leading to inevitable linear and pitch coupling).
Here is a bode plot of the EDC without the bar. It seems very similar to the bode plot with the bar

Here is a bode plot of the rigidly-supported EDC, without bar. I need to do a comparison plot of the rigid and flexibly-supported EDCs (without bar)

|
5194
|
Thu Aug 11 16:07:37 2011 |
steve | Update | SUS | ETMY rack cables strain releived |
|
5193
|
Thu Aug 11 15:31:43 2011 |
steve | Update | General | malfunctioning crane at ETMY |
Quote: |
The horizontal trolley drive stopped working at the east end this morning. It is working intermittently. In the worst case we can take the door off with the manual -Genie- lift.
I'm working with Konecrane to solve the wormgear drive problem.
|
Gear box found, it's lead time one week at $825 The crane may be functional round August 25 |
5192
|
Thu Aug 11 14:32:12 2011 |
Koji | Summary | SUS | Photosensor Head Calibration Curve for TT Frame |
The entry was quite confusing owing to many misleading wordings.
- The PS2 should be calibrated "as is". (i.e. should be calibrated with the frame)
- The previous calibrations with the highly reflective surface were 0.32V/mm and 0.26V/mm, respectively.
This time you have 0.10V/mm (with an undescribed surface). The ratio is not 32 but 3.2.
- The DC output of PS2 on the shaking setup was 2.5V. The DC output seen in the plot is 3.5V-ish.
This suggests the possibiliteies:
1) The surface has slightly higher reflectivity than the frame
2) The estimation of the distance between the frame and the PS2 during the TF measurement was not accurate.
- The word "DC coupling level" is misleading. I guess you mean the DC value of the vbration isolation transfer function
of the suspension.
Quote: |
I have re-calibrated the photosensor I used to measure the displacements of the TT frame (what I call "Photosensor 2").
As before, the linear region is about 15.2mm to 25.4mm. It is characterized by the slope -0.0996 V/mm (-0.1 V/mm). Recall that photosensor 1 (used to measure mirror displacements) has a calibration slope of -3.2V/mm. The ratio of the two slopes (3.2/0.1 = 32). We should thus expect the DC coupling level to be 32? This is not what we have for the DC coupling levels in our data (2.5 for flexibly-supported, fully-assembled TT (with EDC, with bar), 4.2 for EDC without bar, 3.2 for rigid EDC without bar, 3.2 for no EDC, with bar, 3.2 for no EDC without bar) . I think I may need to do my calibration plot for the photosensor at the frame?

|
|
5191
|
Thu Aug 11 14:22:00 2011 |
Nicole | Summary | SUS | Photosensor Head Calibration Curve for TT Frame |
Quote: |
I have re-calibrated the photosensor I used to measure the displacements of the TT frame (what I call "Photosensor 2").
As before, the linear region is about 15.2mm to 25.4mm. It is characterized by the slope -0.0996 V/mm (-0.1 V/mm). Recall that photosensor 1 (used to measure mirror displacements) has a calibration slope of -3.2V/mm. The ratio of the two slopes (3.2/0.1 = 32). We should thus expect the DC coupling level to be 32? This is not what we have for the DC coupling levels in our data (2.5 for flexibly-supported, fully-assembled TT (with EDC, with bar), 4.2 for EDC without bar, 3.2 for rigid EDC without bar, 3.2 for no EDC, with bar, 3.2 for no EDC without bar) . I think I may need to do my calibration plot for the photosensor at the frame?

|
I have redone the voltage versus displacement measurements for calibrating "Photosensor 2" (the photosensor measuring the motions of the TT frame). This time, I calibrated the photosensor in the exact position it was in during the experimental excitation ( with respect to the frame ). I have determined the linear region to be 15.2mm to 22.9mm (in my earlier post today, when I calibrated the photosensor for another location on the frame, I determined the linear region to be 15.2mm to 25.4mm). This time, the slope was -0.92 V/mm (instead of -0.1 V/mm).
This means that the calibration ratio for photosensor 1 (measuring mirror displacements) and photoensor 2 (measuring frame displacements) is 34.86.
Since this "unity" value should be 34.86 for my transfer function magnitude plots (instead of the ~3 value I have), do I need to scale my data? It is strange that it differs by an order of magnitude... |
5190
|
Thu Aug 11 13:41:36 2011 |
Ishwita, Manuel | Update | PEM | Coherence of Guralp1 and STS2(Bacardi, Serial NR 100151) |
Following is the coherence plot obtained when Guralp1 and STS2(Bacardi, Serial NR 100151) are placed very close to each other (but they aren't touching each other):

The seismometers were placed as shown in the picture below:

They are placed below the center of the mode cleaner vacuum tube. |
5189
|
Thu Aug 11 12:54:06 2011 |
kiwamu | Update | IOO | MC spot positions |
The spot positions on the MC mirrors were adjusted by steering the MC mirrors, resulting in 1 mm off-centering on each optic.
DONE.
(Requirement cleared)
One of the requirements in aligning the MC mirrors is the differential spot positions in MC1 and MC3.
It determines the beam angle after the beam exists from MC, and if it's bigger than 3 mm then the beam will be possibly clipped by the Faraday (#4674).
The measured differential spot positions on MC1 and MC3 are : PIT = 0.17 mm and YAW = 1.9 mm, so they are fine.
(Measurement and Results)
Suresh and I aligned the MC cavity's eigen axis by using MCASS and steering the MC mirrors.
Most of the alignment was done manually by changing the DC biases
because we failed to invert the output matrix and hence unable to activate the MCASS servo (#5167).
Then I ran Valera's script to measure the amount of the off-centering (#4355), but it gave me many error messages associated with EPICS.
So a new script newsensedecenter.csh, which is based on tdsavg instead of ezcaread, was made to avoid these error messages.
The resultant plot is attached. The y-axis is calibrated into the amount of the off-centering in mm.
In the plot each curve experiences one bump, which is due to the intentional coil imbalance to calibrate the data from cnts to mm (#4355).
The dashed lines are the estimated amount of off-centering.
For the definition of the signs, I followed Koji's coordinate (#2864) where the UL OSEM is always in minus side.
|
Feb 26 2011 |
May 08 2011 |
Aug 2 2011 |
[NEW!!] Aug 10 2011 (in air) |
MC1 pit [mm] |
1.6 |
1.9 |
1.93 |
-0.858 |
MC2 pit [mm] |
6.4 |
9.0 |
9.03 |
-0.844 |
MC3 pit [mm] |
1.4 |
2.0 |
2.01 |
-1.03 |
MC1 yaw [mm] |
-1.5 |
-1.7 |
-1.72 |
-0.847 |
MC2 yaw [mm] |
1.0 |
0.2 |
0.178 |
0.582 |
MC3 yaw [mm] |
-1.3 |
-1.9 |
-1.87 |
-1.06 |
Quote from #5182 |
After the beam spots on MC1 and MC3 were close to the actuation nodes (<1mm away)
|
|
5188
|
Thu Aug 11 12:31:39 2011 |
Nicole | Summary | SUS | Photosensor Head Calibration Curve for TT Frame |
I have re-calibrated the photosensor I used to measure the displacements of the TT frame (what I call "Photosensor 2").
As before, the linear region is about 15.2mm to 25.4mm. It is characterized by the slope -0.0996 V/mm (-0.1 V/mm). Recall that photosensor 1 (used to measure mirror displacements) has a calibration slope of -3.2V/mm. The ratio of the two slopes (3.2/0.1 = 32). We should thus expect the DC coupling level to be 32? This is not what we have for the DC coupling levels in our data (2.5 for flexibly-supported, fully-assembled TT (with EDC, with bar), 4.2 for EDC without bar, 3.2 for rigid EDC without bar, 3.2 for no EDC, with bar, 3.2 for no EDC without bar) . I think I may need to do my calibration plot for the photosensor at the frame?

|
5187
|
Thu Aug 11 11:50:56 2011 |
Manuel , Ishwita | Update | PEM | Calibration of Guralp and STS2 |
We just checked with a function generator the calibration of the ADC. We set a square wave with amplitude 1V. We measured the voltage with the oscilloscope and we found on the data viewer that one volt is 3208 counts. That's what we expected (+/- 10V for 16bits) but now we are more sure.
|
5186
|
Thu Aug 11 10:56:08 2011 |
Ishwita, Manuel | Update | PEM | Moving Seismometers |
Quote: |
We turned off the power of the seismometers and moved the Guralp1 close to the STS. Both are now situated below the center of the mode cleaner vacuum tube.
We oriented the X axis of the STS & Guralp1 along the X axis of the interferometer. Then we turned on the power again, but the STS channels don't give any signal. We think this is, because we didn't push the auto zero button.
|
After pressing the auto-zero button (a lot of times) of the STS breakout box & aligning the bubble in the STS, we could finally get data from STS (Bacardi). So, now STS2 (Bacardi - Serial NR. 100151) is working! |
5185
|
Thu Aug 11 09:39:25 2011 |
Ishwita, Manuel | Update | PEM | Calibration of Guralp and STS2 |
Quote: |
Quote: |
I'm pretty sure that don't have any ADC's with this gain. It should be +/- 10V for 16 bits.
|
Jenne told us that the ADC was +/- 2V for 16 bits so our calibration is wrong. Since, the ADC is +/- 10V for 16 bits we need to change our calibration and now we can also use the purple STS breakout box.
|
New calibration for Guralp:
ADC: 216counts = 20V Hence, calibration of ADC is (215x0.1) counts/V.
GURALP
Sensitivity = 800 V/ms-1
(215 x 0.1) counts/V x 800 V/ms-1 = 2621440 counts/ms-1 -----> 3.8147e-07 ms-1/count
Calibration = 3.8147e-07 ms-1/count
Using the above calibration we obtain the following plot:

When we compare this plot with the old plot (see here) we see that in our calibration, we have got a factor of 10 less than the old plot. We do not know the gain of the Guralp. If we assume this missing 10 factor to be the gain of Guralp then we can get the same calibration as the old plot. But is it correct to do so? |
5184
|
Thu Aug 11 08:29:28 2011 |
steve | Update | Computers | dataviewer at Rosalba |
I'm having this problem with DTV every morning at Rosalba only. It wants to start with a negative GPS time and it can not connect to the frame builder.
Normally after a few time of starting it, it will work. |
5183
|
Thu Aug 11 06:45:14 2011 |
Nicole | Summary | SUS | Shaking Testing |
Koji and I have finished shaking the table for the first round of measurements (horizontal shaking). We have cleaned up the lab space used.
The FFT Analyzer has been put back to its position at the back side of the rack (near the seismometers).
I will calibrate the photosensor for the suspension frame and piece together/analyze/produce graphs of the data today. If everything is fine (the measurements are fine) and if there is a chance, we hope to shake the TT suspension vertically. |
5182
|
Thu Aug 11 04:45:07 2011 |
Suresh | Update | IOO | Aligning the 1064nm beam with the in-vacuum pzt's |
[Kiwamu, Suresh]
We worked on the beam path from MC to BS this evening.
After the beam spots on MC1 and MC3 were close to the actuation nodes (<1mm away) we checked the beam position on the Faraday Isolator (FI) to make sure that it is passing through both the input and output apertures without clipping. The beam is slightly displaced (by about half a beam diameter) downwards at the input of the FI. The picture below is a screen shot from the MC1 monitor while Kiwamu held an IR card in front of the FI.

We then proceeded to check the beam position on various optical elements downstream. But first we levelled the BS table and checked to see if the reflection from PJ1 (1st Piezo) is landing on the MMT1 properly. It was and we did not make any adjustment to PJ1. However the reflection from MMT1 was not centered on MMT2. We adjusted the MMT1 to center the beam on it. We then adjusted MMT2 to center the beam on PJ2. At this point we noticed that the spot on IPPO (pick off window) was off towards the right edge. When we centered the beam on this it missed the center of the PRM. In order to decide what needs to be moved, we adjusted PJ2 such that the beam hits the PR2, bounces back to PR3, and becomes co-incident with the green beam from X-arm on the BS. Under this condition the beam is not in the center of PRM and nor is it centered on IPPO. In fact it is being clipped at the edge of the IPPO.
It is clear that both IPPO and the PRM need to be moved. To be sure that the beam is centered on PR2 we plan to open the ITMX chamber tomorrow. |
5181
|
Thu Aug 11 02:16:57 2011 |
Jenne | Update | Green Locking | Y-green aligned and flashing |
[Jenne, with ample supervision by Kiwamu and Suresh]
Y-green was aligned, and is now flashing. The ETMY trans camera was very helpful for this alignment. I didn't end up needing to use a foil aperture.
Kiwamu and Suresh had just closed up the IOO doors, so we don't know yet where it's hitting on the PSL table (if the beam is making it that far). Tomorrow we'll look at ITMY to see if the green beam is centered there, and if it's coming out to the PSL table. |
5180
|
Wed Aug 10 22:47:22 2011 |
rana | Summary | VAC | Vacuum Workstation (linux3) re-activated |
For some reason the workstation at the vac rack was off and unplugged. Nicole and I plugged its power back in to the EX rack.
I turned it on and it booted up fine; its not dead. To get it on to the network I just made the conversion from 131.215 to 192.168 that Joe had done on all the other computers several months ago.
Now it is showing the Vacuum overview screen correctly again and so Steve no longer has to monopolize one of the Martian laptops over there. |
5179
|
Wed Aug 10 20:40:17 2011 |
Jenny | Update | PSL | PDH locking: got an error signal |
I ended up choosing a different dither frequency for driving the NPRO PZT: 230 kHz, because the phase modulation response in that region is higher according to other data taken on an NPRO laser (see this entry). At 230 there is a dip in the AM response of the PZT.
I am driving the PZT at 230 kHz and 13 dBm using a function generator. I am then monitoring the RF output of a PD that is detecting light reflected off the cavity. (The dither frequency was below the RF cutoff frequency of the PD, but it was appearing in the "DC output", so I am actually taking the "DC output" of the PD, which has my RF signal in it, blocking the real DC part of it with a DC block, and then mixing the signal with the 230kHz sine wave being sent to the PZT.
I am monitoring the mixer output on an oscilloscope, as well as the transmission through the cavity. I am sweeping the laser temperature using a lock in as a function generator sending out a sine wave at 0.2 V and 5 mHz. When there is a peak in the transmission, the error signal coming from the mixer passes through zero.
My next step is to find or build a low pass filter with a pole somewhere less than 100 kHz to cut out the unwanted higher frequency signal so that I have a demodulated error signal that I can use to lock the laser to the cavity.
|
5178
|
Wed Aug 10 19:18:26 2011 |
Nicole | Summary | SUS | Fixed Reflective Photosensors; Recalibrated Photosensor 2 |
Thanks to Koji's help, the second photosensor, which was not working, has been fixed. I have re-calibrated the photosensor after fixing a problem with the circuit. I have determined the new linear region to lie between 7.6 mm and 19.8mm. The slope defining the linear region is -0.26 V/mm (no longer the same as the first photosensor, which is -0.32 V/mm).
Here is the calibration plot.

|